OpenStudy (anonymous):

Find all solutions to the equation. (sin x)(cos x) = 0

OpenStudy (jdoe0001):

well, that equation will only be 0 when either one is 0 so cosine is 0 at the 0 and \(\pi \) and \(2 \pi\) and sine is 0 at \(\cfrac{\pi}{2}, \cfrac{3\pi}{2}\)

OpenStudy (anonymous):

Simply find all places where either the sin x or the cos x is "0". That is:\[\frac{ k \pi }{ 2 }\]where k is any integer

OpenStudy (jdoe0001):

where either is 0 or both are 0, at those angles, your equation is = 0 so just rewrite in a cyclical notation, like tcarroll010 above

OpenStudy (anonymous):

my answers are kπ, (π/2)+ kπ, kπ (π/2) + 2kπ or (π/2) + kp

OpenStudy (anonymous):

To fit the format I gave, you would go with choice #2.

OpenStudy (anonymous):

Choice #2 is the same as my answer, but I do like the way I wrote it because of the neat compactness.

OpenStudy (anonymous):

All good now, @laxman03 ?

OpenStudy (anonymous):

yes thankyou very much.

OpenStudy (anonymous):

uw!

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @laxman03